Toss a 4 coins 15 times and write your outcomes, what will be the probability of getting
At least 3 head
At most 2 head
Exactly 1 head

Answer by ewatrrr(24785)


Hi

Binomial Distribution:  p = 1/2 and q = 1/2  n = 4 (coins)
Using Calculator (TI instructions given here)
An inexpensive, easy to use  Casio fx-115 ES plus works similarly

P(At least 3 head)= P(x ≥ 3) =   1- P(x ≤ 2) 
        1 - binomcdf(4,.5,2) = 1 - .6875 = .3125  0r 31.25%

P(At most 2 head) = P(x ≤ 2) = binomcdf(4,.5,2) =.6875  0r 68.75%

P(Exactly 1 head) = P(x=1) = binompdf(4,.5,1) = .25 0r 25%

One would expect, over the 15 times of tossing the 4 coins..
on an average... the above percentages would hold true.

Wish You the Best in your Studies.